tickets for a movie cost $7.25 for adults and $5.75 for children. One day 12 tickets were sold for $81.00. How many of each ticket were sold?
Answer by algebrahouse.com(1659)

x = number of children tickets sold
12 - x = number of adult tickets sold {total number of tickets sold was 12}

5.75x + 7.25(12 - x) = 81 {price of ticket times number of tickets equals total price}
5.75x + 87 - 7.25x = 81 {used distributive property}
-1.5x + 87 = 81 {combined like terms}
-1.5x = -6 {subtracted 87 from each side}
x = 4 children tickets {divided each side by -1.5}
12 - x = 8 adult tickets {substituted 4, in for x, into 12 - x}

4 children tickets
8 adult tickets

For more help from me, visit: www.algebrahouse.com